According to 2016 Census data, the following probability model shows the different modes of transportation Americans...

  1. According to 2016 Census data, the following probability model shows the different modes of transportation Americans used to get to work on a given day.

Mode

Reg. Freg

Drove Alone

76.3%

Carpooled

9%

Public Transit

5.1%

Cycled

0.6%

Walked

2.8%

Taxi

1.2%

Worked from Home

5%

  1. Explain why this would be considered a probability model.
  2. If an adult is randomly selected in the US, what is the probability that they will have drove alone or walked to work that day?
  3. What is the probability that a randomly selected person did not carpool to work today?
Answer #1

a) sum of all probability = 0.763 + 0.09 + 0.051 + 0.006 + 0.028 + 0.012 + 0.05 = 1

b) P(drove alone or walked to work) = P(drove alone) + P(walked) = 0.763 + 0.028 = 0.791

c) P(not carpool) = 1 - P(carpool) = 1 - 0.09 = 0.91